OpenStudy (anonymous):

how do i use the distributive law to factor 2x+22+18y

Parth (parthkohli):

Factor the \(2\) out

OpenStudy (whpalmer4):

\[2x + 22 + 18 y = 2*1x + 2*11 + 2*9y = 2(something with three terms)\]

OpenStudy (anonymous):

so will it be 2(x+11+9y?

OpenStudy (whpalmer4):

Yes, \[2(x+11+9y)\]